A 500g uniform sphere of 7cm radius spins at 30rev/sec on an axis through its center find its
A, KEr
B, angular momentum
C, radius of gyration
1
Expert's answer
2020-05-12T09:31:06-0400

The moment of inertia of a solid uniform sphere is


"I=\frac{2}{5}mr^2."

The kinetic energy of rotation is


"KE=\frac{1}{2}I\omega^2=\frac{1}{5}mr^2\omega^2=\frac{1}{5}mr^2(2\pi n)^2=\\
=17.4\text{ J}."

The angular momentum is


"L=I\omega=\frac{4\pi}{5}mnr^2=0.185\text{ Js}."

The radius of gyration is


"R_g=\sqrt\frac{I}{m}=r\sqrt\frac{2}{5}=4.43\text{ cm}."